# Anisotropic linear waves and breakdown of the momentum expansion in spin magnetohydrodynamics

### v1

@@ A.2 Eigenmodes at $k_{\parallel}=0$
00 0 $=$ $\left(\begin{array}[]{ccc|ccc}-i\frac{c_{s}^{2}\rho_{\perp}^{\prime}}{1-v_{A}^{2}}\frac{B}{h}k_{\perp}^{2}&-k_{\perp}B&\omega+i\rho^{\prime}_{\perp}k_{\perp}^{2}&&\\ \omega&-hk_{\perp}&\frac{h}{B}v_{A}^{2}\omega&&\\ -c_{s}^{2}k_{\perp}&h\omega+i(\zeta_{\perp}+\eta_{\perp})k_{\perp}^{2}&-\frac{h}{B}v_{A}^{2}k_{\perp}&\\ \hline\cr&&&h(1-v_{A}^{2})\omega+i(\eta_{\parallel}+2\xi_{\parallel}+\gamma_{\parallel})k_{\perp}^{2}&&-\frac{4}{\chi}(\xi_{\parallel}+\gamma_{\parallel})k_{\perp}\\ &&&2(\gamma_{\parallel}+\xi_{\parallel})k_{\perp}&&\omega+i\Gamma_{\parallel}\end{array}\right)\begin{pmatrix}\delta\epsilon\\ \delta u_{x}\\ \delta B_{z}\\ \delta u_{z}\\ \delta\sigma_{y}\end{pmatrix},$ (65f)
00 0 $=$ $\left(\begin{array}[]{cc|cc}h\omega+i(\eta_{\perp}+\gamma_{\perp})k_{\perp}^{2}&\frac{4}{\chi}\gamma_{\perp}k_{\perp}&&\\ -2\gamma_{\perp}k_{\perp}&\omega+i\Gamma_{\perp}&&\\ \hline\cr&&\omega+i\rho_{\parallel}^{\prime}k_{\perp}^{2}&\\ &&&\omega+i\Gamma_{\parallel}\end{array}\right)\begin{pmatrix}\delta u_{y}\\ \delta\sigma_{z}\\ \delta B_{y}\\ \delta\sigma_{x}\end{pmatrix}.$ (65l)
@@
±$\pm v_{f}k_{\perp}-\frac{i}{2}\Big{(}\frac{v_{A}^{2}}{v_{f}^{2}}(1-c_{s}^{2})^{2}\rho^{\prime}_{\perp}+\frac{\eta_{\perp}+\zeta_{\perp}}{h}\Big{)}k_{\perp}^{2},$
@@ Appendix B Explicit forms of the cosine expansion
×$\times(\omega-\tilde{\omega}_{3}-\omega_{3}\cos^{2}\theta)(\omega-\tilde{\omega}_{4}-\omega_{4}\cos^{2}\theta).$
@@
As already discussed, the ${\mathcal{O}}(k^{2})$ terms are out of our interest as they are beyond the first-order spin MHD. Expanding Eq. (71) up to the $k^{2}$ order, one gets the parameters $\Omega_{i}$ for Eq. (59) as
@@
$+$ $h^{3}\left(1-{v_{A}}^{2}\right){v_{f}}^{2}\chi\left((\gamma_{\parallel}+\xi_{\parallel})^{2}-\gamma_{\parallel}\tilde{\eta}_{\parallel}\right)\Big{)}$
@@
×$\times\Big{[}{v_{A}}^{2}\Big{(}2\gamma_{\parallel}\Big{(}\left(1-{v_{A}}^{2}\right)\frac{\left(1-{c_{s}}^{2}\right)^{2}h{\rho_{\perp}}{v_{A}}^{2}}{{v_{f}}^{2}}$
@@
-- $h{\rho_{\perp}}\Big{(}\frac{8\gamma_{\parallel}{c_{s}}^{4}h{\rho_{\perp}}{v_{A}}^{2}}{{v_{f}}^{2}}-8{v_{A}}^{2}\gamma_{\parallel}(\zeta_{\parallel}-\tilde{\eta}_{\parallel})$

### v2

@@ A.2 Eigenmodes at $k_{\parallel}=0$
$0$ $=$ $\left(\begin{array}[]{ccc|ccc}-i\frac{c_{s}^{2}\rho_{\perp}^{\prime}}{1-v_{A}^{2}}\frac{B}{h}k_{\perp}^{2}&-k_{\perp}B&\omega+i\rho^{\prime}_{\perp}k_{\perp}^{2}&&\\ \omega&-hk_{\perp}&\frac{h}{B}v_{A}^{2}\omega&&\\ -c_{s}^{2}k_{\perp}&h\omega+i(\zeta_{\perp}+\eta_{\perp})k_{\perp}^{2}&-\frac{h}{B}v_{A}^{2}k_{\perp}&\\ \hline\cr&&&h(1-v_{A}^{2})\omega+i(\eta_{\parallel}+2\xi_{\parallel}+\gamma_{\parallel})k_{\perp}^{2}&&-\frac{4}{\chi}(\xi_{\parallel}+\gamma_{\parallel})k_{\perp}\\ &&&2(\gamma_{\parallel}+\xi_{\parallel})k_{\perp}&&\omega+i\Gamma_{\parallel}\end{array}\right)\begin{pmatrix}\delta\epsilon\\ \delta u_{x}\\ \delta B_{z}\\ \delta u_{z}\\ \delta\sigma_{y}\end{pmatrix}$, (71f)
$0$ $=$ $\left(\begin{array}[]{cc|cc}h\omega+i(\eta_{\perp}+\gamma_{\perp})k_{\perp}^{2}&\frac{4}{\chi}\gamma_{\perp}k_{\perp}&&\\ -2\gamma_{\perp}k_{\perp}&\omega+i\Gamma_{\perp}&&\\ \hline\cr&&\omega+i\rho_{\parallel}^{\prime}k_{\perp}^{2}&\\ &&&\omega+i\Gamma_{\parallel}\end{array}\right)\begin{pmatrix}\delta u_{y}\\ \delta\sigma_{z}\\ \delta B_{y}\\ \delta\sigma_{x}\end{pmatrix}$. (71l)
@@
$\pm v_{f}k_{\perp}-\frac{i}{2}\Big(\frac{v_{A}^{2}}{v_{f}^{2}}(1-c_{s}^{2})^{2}\rho^{\prime}_{\perp}+\frac{\eta_{\perp}+\zeta_{\perp}}{h}\Big)k_{\perp}^{2},$
@@ Appendix B Explicit forms of the cosine expansion
$\times(\omega-\tilde{\omega}_{3}-\omega_{3}\cos^{2}\theta)(\omega-\tilde{\omega}_{4}-\omega_{4}\cos^{2}\theta).$
@@
As already discussed, the ${\mathcal{O}}(k^{2})$ terms are outside of our interest, as they are beyond the first-order spin MHD. Expanding Eq. (77) up to the $k^{2}$ order, one gets the parameters $\Omega_{i}$ for Eq. (61) as
@@
$+$ $)\displaystyle h^{3}\left(1-{v_{A}}^{2}\right){v_{f}}^{2}\chi\left((\gamma_{\parallel}+\xi_{\parallel})^{2}-\gamma_{\parallel}\tilde{\eta}_{\parallel}\right$$\Big)$
@@
$\times\Big[{v_{A}}^{2}\Big(2\gamma_{\parallel}\Big(\left(1-{v_{A}}^{2}\right)\frac{\left(1-{c_{s}}^{2}\right)^{2}h{\rho_{\perp}}{v_{A}}^{2}}{{v_{f}}^{2}}$
@@
[…]